OpenStudy (anonymous):

Can you put this in standard form please? y = 1(x-11)^2 + 30

OpenStudy (mrnood):

you mean y = ax^2 +bx + c ?

OpenStudy (mrnood):

\[y=(x-11)^{2}+30\] Is that the equation you start with (you have an extra 1 in your first post...)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

?

OpenStudy (anonymous):

y = 2x + 8*

OpenStudy (mrnood):

so take the bracket \[(x-11)^{2}\] and expand it

OpenStudy (anonymous):

y = 2x - 22, which simplifies to y = 2x + 8 if you include the 30 correct?

OpenStudy (mrnood):

NO the bracket is 'squared' - you are simply multiplying it by 2

OpenStudy (anonymous):

So it'd be x^2 + 121? since the 2 negatives cancel

OpenStudy (mrnood):

slow down do oyu know how to expand (a+b)^2 ?

OpenStudy (anonymous):

No, isn't it (a+b)(a+b)?

OpenStudy (anonymous):

and you use the distributive property?

OpenStudy (mrnood):

That is correct - but that is not what you wrote above

OpenStudy (anonymous):

Oh wow, Idk what I just did lol so it'd be y = (x-11)(x-11) + 30 ?

OpenStudy (mrnood):

Yes

OpenStudy (anonymous):

y = 2x -11x -11x + 121 + 30?

OpenStudy (mrnood):

NO You are still multiplying by 2 rather than squaring

OpenStudy (mrnood):

your last post was close - but 1 mistake

OpenStudy (anonymous):

oh x^2?

OpenStudy (mrnood):

that's it - so complete the sum - group all the factors together ... and voila - standard form

OpenStudy (anonymous):

y = x^2 -22x + 151?

OpenStudy (mrnood):

That's correct - just check your original though - why did oyu write a 1 in front of the bracket?

OpenStudy (anonymous):

Wow thanks. Um, that's what A was substituted for in the original equation, she wants me to show all work, even the 1.

OpenStudy (anonymous):

Appreciate your help!

OpenStudy (mrnood):

NP - outta here now